ELECTRONIC GROUND AND EXCITED STATES OF CoN: AN AB INITIO MOLECULAR ORBITAL STUDY

\underline{TSUNEO HIRANO} and UMPEI NAGASHIMA, Grid Technology Research Center, Institute of Advanced Industrial Science and Technology, 6-9-3 Ueno, Taito-ku, Tokyo 110-0015, Japan.

Based on the information from Tanaka group of Hokkaido University,^{a} we have studied the ^{1}Σ^{+} and ^{5}Δ states of CoN by ab initio molecular orbital methods as an extension of our previous work on FeN.^{b} The MR-SDCI+Q+E_{rel}/[Roos ANO(Co), aug-cc-pVQZ(N)] calculations^{c} with full-valence plus Co 3s and 3p electron correlations predicted that the ^{5}Δ level should be located by about 4300 cm^{−1} higher than the ^{1}Σ^{+} state. Hence, the electronic ground state is ^{1}Σ^{+}. The equilibrium bond lengths for the ^{1}Σ^{+} and ^{5}Δ states at this level of calculation are 1.5621 and 1.5945 Å, respectively. The first order relativistic correction E_{rel} increases linearly with the Co-N bond length, with steeper gradient for ^{1}Σ^{+} than 5Δ states as is expected. The resultant shortening of the Co-N bond due to the relativistic effect is 0.016 and 0.005 Å for the ^{1}Σ^{+} and ^{5}Δ states, respectively.

aK. Tanaka and M. Sekiya, Division of Chemistry, Graduate School of Science, Hokkaido University, Sapporo 060-0808, Japan.

bM. Amano, S.S Itono, T. Hirano, U. Nagashima, M. Sekiya, and K. Tanaka, The 57th Ohio State University International Symposium on Molecular Spectroscopy, FB06, 255 (2002)

cQ: Davidson’s correction; E_{rel}: Relativistic correction